OpenStudy (anonymous):

what is product (3x^2 + 2x - 3)(x - 1)

OpenStudy (ivettef365):

multiply as follows: (3x^2)(x) = 3x^3 (3x^2)(-1) = -3x^2 (2x)(x) = 2x^2 (2x)(-1) = -2x you can do the last one with (-3)

OpenStudy (anonymous):

(3x^2 + 2x - 3)(x - 1) = x(3x^2 + 2x - 3) -1(3x^2 + 2x - 3) =3x^3+2x^2-3x-3x^2 - 2x +3 =3x^3+(2x^2-3x^2) -3x - 2x +3 =3x^3+(-x^2) -5x +3 =3x^3-x^2 -5x +3 Therefore (3x^2 + 2x - 3)(x - 1)=3x^3-x^2 -5x +3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome